George leaves his house on his bike at 7:30 AM and averages 5 miles per hour. His wife leaves at 8 AM following the same path averaging 8 miles per hour. At what time will his wife catch George?
Answer by mananth(16949)

George speed 5 mph
Wife speed 8 mph
George speed 07:30
Wife speed 08:00
Difference in time= 00:30 => 0.5 hours
George speed will have covered 2.5 miles before wife starts
catch up distance= 2.5 miles
catch up speed = 8 -5 mph
catch up speed = 3 mph
Catchup time = catchup distance/catch up speed
catch up time= 2.5 / 3
catch up time= 5/6 hours =50 minutes
she will catch up at 8:50 AM